# PoGDiff: Product-of-Gaussians Diffusion Models for Imbalanced Text-to-Image Generation

### v2

@@ 4.2 Results
Generation Quality and Accuracy. We report the performance of different methods in terms of FID score, human evaluation score, GPT-4o score, and DINO score in Table 1,  Table 2,  Table 3 and  Table 4, respectively [^2] CLIP score is not applicable here. Our text prompts are predominantly human names while CLIP is primarily trained on common objects, not human names; therefore CLIP score cannot measure the matching between images and human names. . Across all tables, we observe that our PoGDiff consistently outperforms all baselines. Notably, PoGDiff demonstrates significant improvements, especially in few-shot scenarios (i.e., for minority classes). It is also worth noting that CBDM [26] performs extremely poorly on AgeDB-IT2I-S and AgeDB-IT2I-M datasets. This is because their method samples text conditions from the entire space, which may work in one-hot class settings, but in our context (natural text conditions), this sampling technique misguides the model during training.
Fig. 5 shows randomly sampled generated images on low-density classes (Column $1\&2$) and high-density class (Column $3$) in AgeDB-IT2I-L for each method. Note that the ground-truth (GT) images are the training images. For the high-density class, we select $10$ out of $24$ total images in training set to report in this figure. Across each column, the individual names are Albert Einstein, JW Marriott, and Luise Rainer, respectively. PoGDiff achieves significantly better accuracy and quality for both head and tail classes (see Appendix G for more comparisons and analysis). Specifically, both SDv1.5 and T2H fail to generate accurate individuals (green boxes), and CBDM even struggles to generate images with correct genders (blue boxes). By contrast, our PoGDiff successfully generates accurate individuals, even when trained on a dataset containing only two images.
Datasets AgeDB-IT2I VGGFace-IT2I
Size Small Medium Large Large
Metric gRecall Score $\uparrow$
Shot All Few All Few All Few All Few
Vanilla 0.017 0.000 0.104 0.167 0.196 0.200 0.133 0.167
CBDM 0.267 0.000 0.159 0.083 0.138 0.100 0.120 0.100
T2H 0.017 0.000 0.104 0.167 0.196 0.200 0.133 0.167
PoGDiff (Ours) 0.800 1.000 0.517 0.642 0.435 0.540 0.400 0.533
Table 6: Ablation studies for PoGDiff.
Datasets AgeDB-IT2I-Large
Size FID $\downarrow$ Human  $\uparrow$ GPT-4o $\uparrow$ DINO $\uparrow$
Shot All Few All Few All Few All Few
w/o ${\bf y}^{\prime}$ (Vanilla) 7.67 11.67 0.60 0.20 4.90 3.60 0.34 0.25
w/o $\psi_{\textrm{img-sim}}$ 6.41 10.49 0.84 0.68 8.40 7.60 0.57 0.46
w/o $\psi_{\textrm{inv-txt-den}}$ 6.35 10.43 0.84 0.68 8.20 7.80 0.64 0.51
PoGDiff (Ours) 6.03 10.16 0.84 0.68 8.50 8.00 0.66 0.52
To see why, note that the FID score measures the distance between two Gaussian distributions, where the mean of the Gaussian represents the identity (ID) and the variance represents the diversity. For example, the mean of the ground-truth distribution represents the embedding position of the ground-truth ID, while the variance of the ground-truth distribution represents the diversity of ground-truth images, and similarly for the generated images. Therefore, a lower FID score indicates that the generated-image distribution better matches the ground-truth distribution in terms of both ID and diversity.
Discussion on Generation Diversity. Table 1 and Fig. 5 demonstrate our PoGDiff’s promising generation diversity:
@@
Results on gRecall. Table 5 shows the gRecall for different methods on three datasets, AgeDB-IT2I-small, AgeDB-IT2I-medium, and AgeDB-IT2I-large. These results show that our PoGDiff achieves much higher gRecall compared to all baselines, demonstrating its impressive diversity and coverage of different attributes of the same individual in the training set (see Appendix F for more discussion and examples on gRecall).
Ablation Study. To verify the effectiveness of each component in the second term in our PoGDiff final objective function from Eqn. (12), we report the accuracy of our proposed PoGDiff after removing the ${\bf y}^{\prime}$ (i.e., same as Vanilla model), the Image Similarity term $\psi_{\textrm{img-sim}}$, and/or the Inverse Text Densities term $\psi_{\textrm{inv-txt-den}}$ in Table 6 for AgeDB-IT2I-L. The results show that removing either term leads to a performance drop, confirming the importance of both terms in our PoGDiff.
@@ 5 Conclusions
In this paper, we propose a general fine-tuning approach called PoGDiff to address the performance drop that occurs when fine-tuning on imbalanced datasets. Instead of directly minimizing the KL divergence between the predicted and ground-truth distributions, PoGDiff replaces the ground-truth distribution with a Product of Gaussians (PoG), constructed by combining the original ground-truth targets with the predicted distribution conditioned on a neighboring text embedding. Looking ahead, an interesting avenue for future research would be to explore more innovative techniques for re-weighting minority classes (as discussed in Sec. E), particularly within the constraints of: (1) long-tailed generation settings, as opposed to recognition tasks, and (2) natural text prompts rather than one-hot class labels. Exploring PoGDiff for other modalities (e.g., videos and time series) is also an interesting future work.
Impact Statement
@@
Appendix B Details for Baselines
@@
  • T2H: We use the term T2H to denote a model that uses Long-Tailed Diffusion Models with Oriented Calibration (T2H) [40]. T2H is a reweighting method similar to CBDM [26], but is not directly applicable to our setting. Specifically, T2H [40] relies on the class frequency, which is not available in our setting. In this paper, we adapted this method to our settings by using the density for each text prompt embedding to serve as the class frequency in T2H [40].
Appendix C Details for Evaluation
@@
FID Score. For each identity, we collect all images from the original AgeDB or DigiFace datasets as the true image set. Then, In all-shot evaluation, for AgeDB-IT2I-S and AgeDB-IT2I-M, we generate $100$ images per identity as the fake image set, and for AgeDB-IT2I-L and DigiFace-IT2I-L, we generate $20$ images per identity as the fake image set. In few-shot evaluation, we we generate $500$ images per identity as the fake image set. For all generations, we employ the DDIM sampling technique [34] with $50$ steps. The prompt used during generation is “An image of { p }.” where “p” is the name of the identity (e.g., Albert Einstein).
@@
Appendix D Details for Implementation
@@
Appendix E Limitations
@@
Appendix F Additional Details for AgeDB-IT2I-small in Table 5
@@
Appendix G Discussion
G.1 Problem Settings
@@
Figure 7: Overview of label distribution for four IT2I datasets in bar plots. The x-axis corresponds to the identities (i.e., people or individuals).
Figure 8: TSNE visualization for all the methods for an example individual in the Age-DB-IT2M-large dataset.
G.2 More Detailed Analysis: Understanding Fig. 6
@@
G.3 Why Not Directly Smooth Text Embedding?
@@
G.4 Our Dataset Covers Different Levels of Sparsity

### v3

@@ 4.2 Results
Generation Quality and Accuracy. We report the performance of different methods in terms of FID score, human evaluation score, GPT-4o score, and DINO score in Table 2,  Table 2,  Table 4 and  Table 4, respectively [^2] CLIP score is not applicable here. Our text prompts are predominantly human names, while CLIP is primarily trained on common objects, not human names; therefore, CLIP score cannot measure the matching between images and human names. . Full results are included in Appendix G. Across all tables, we observe that our PoGDiff consistently outperforms all baselines. More results and discussions are in  Appendix J.5. Notably, PoGDiff demonstrates significant improvements, especially in few-shot scenarios (i.e., for minority classes). It is also worth noting that CBDM [15] performs extremely poorly on AgeDB-IT2I-S dataset. This is because their method samples text conditions from the entire space, which may work in one-hot class settings, but in our context (natural text conditions), this sampling technique misguides the model during training.
Fig. 4 shows randomly sampled generated images on low-density classes (Column $1\&2$) and high-density class (Column $3$) in AgeDB-IT2I-L for each method. Note that the ground-truth (GT) images are the training images. For the high-density class, we select $8$ out of $24$ total images in the training set to report in this figure. Across each column, the individual names are Albert Einstein, JW Marriott, and Luise Rainer, respectively. PoGDiff achieves significantly better accuracy and quality for both head and tail classes (see Appendix J for more comparisons and analysis). Specifically, both SDv1.5 and T2H fail to generate accurate individuals (green boxes), and CBDM even struggles to generate images with correct genders (blue boxes). By contrast, our PoGDiff successfully generates accurate individuals, even when trained on a dataset containing only two images.
Generation Diversity. Table 2 and Fig. 4 demonstrate our PoGDiff’s promising generation diversity:
@@
Datasets (-IT2I) AgeDB VGGFace CIFAR-100
Size Small Large Large Large
Metric gRecall Score $\uparrow$
Shot All Few All Few All Few All Few
Vanilla 0.017 0.000 0.196 0.200 0.133 0.167 0.200 0.160
CBDM 0.267 0.000 0.138 0.100 0.120 0.100 0.086 0.067
T2H 0.017 0.000 0.196 0.200 0.133 0.167 0.200 0.160
PoGDiff (Ours) 0.800 1.000 0.435 0.540 0.400 0.533 0.433 0.567
Results on gRecall. Table 5 shows the gRecall for different methods on four datasets, AgeDB-IT2I-small, AgeDB-IT2I-large, VGGFace-IT2I, and CIFAR-100-IT2I. These results show that our PoGDiff achieves much higher gRecall compared to all baselines, demonstrating its impressive diversity and coverage of different attributes of the same individual in the training set (see Appendix I for more discussion and examples on gRecall).
@@ 5 Conclusions
In this paper, we propose a general fine-tuning approach called PoGDiff to address the performance drop that occurs when fine-tuning on imbalanced datasets. Instead of directly minimizing the KL divergence between the predicted and ground-truth distributions, PoGDiff replaces the ground-truth distribution with a Product of Gaussians (PoG), constructed by combining the original ground-truth targets with the predicted distribution conditioned on a neighboring text embedding. Looking ahead, an interesting avenue for future research would be to explore more innovative techniques for re-weighting minority classes (as limitations discussed in Sec. H), particularly within the constraints of (1) long-tailed generation settings, as opposed to recognition tasks, and (2) natural text prompts rather than one-hot class labels. Exploring PoGDiff for other modalities (e.g., videos and time series) is also an interesting future work.
@@
Appendix B Ablation Study
Table 6: Ablation Studies.
Datasets AgeDB-IT2I-Large
Size FID $\downarrow$ Human  $\uparrow$ GPT-4o $\uparrow$ DINO $\uparrow$
Shot All Few All Few All Few All Few
w/o ${\bf y}^{\prime}$ (Vanilla) 7.67 11.67 0.60 0.20 4.90 3.60 0.34 0.25
w/o $\psi_{\textrm{img-sim}}$ 6.41 10.49 0.84 0.68 8.40 7.60 0.57 0.46
w/o $\psi_{\textrm{inv-txt-den}}$ 6.35 10.43 0.84 0.68 8.20 7.80 0.64 0.51
PoGDiff (Ours) 6.03 10.16 0.84 0.68 8.50 8.00 0.66 0.52
To verify the effectiveness of each component in the second term in our PoGDiff final objective function from Eqn. (12), we report the accuracy of our proposed PoGDiff after removing the ${\bf y}^{\prime}$ (i.e., same as Vanilla model), the Image Similarity term $\psi_{\textrm{img-sim}}$, and/or the Inverse Text Densities term $\psi_{\textrm{inv-txt-den}}$ in Table 6 for AgeDB-IT2I-L. The results show that removing either term leads to a performance drop, confirming the importance of both terms in our PoGDiff.
Appendix C Details for Datasets
Note that our method is designed for fine-tuning. Therefore our setup does not require large-scale, long-tailed human datasets. Instead, we sample from these datasets, as long as they meet the following criteria: (1) the dataset must be long-tailed, (2) traditional methods must fail to recognize the minority classes, and (3) there must be a distinguishable difference between the majority and minority classes (e.g., we prefer visual distinctions between the two groups to better highlight the impact of our method). Fig. 6 shows the label density distribution of these datasets, and their level of imbalance (see Appendix J.4 for details on data sparsity).
DigiFace-IT2I: DigiFace-IT2I is derived from the DigiFace dataset [36]. It contains $985$ images across $179$ identities, where each majority class consists of $30$ images and each minority class consists of $2$ images. We use a process similar to AgeDB-IT2I to collect text-image pairs, forming this DigiFace-IT2I dataset.
VGGFace-IT2I: VGGFace-IT2I is a subset from VGGFace2 [37]. It contains $1933$ images across $193$ identities, where each majority class consists of $49$ images and each minority class consists of $2$ images.
Appendix D Details for Baselines
@@
  • T2H: We use the term T2H to denote a model that uses Long-Tailed Diffusion Models with Oriented Calibration (T2H) [30]. T2H is a feature augmentation method, but is not directly applicable to our setting. Specifically, T2H [30] relies on the class frequency, which is not available in our experiments. In this paper, we adapt this method to our settings by using the density for each text prompt embedding to serve as the class frequency in T2H [30].
Appendix E Details for Evaluation
@@
FID Score. For each identity, we collect all images from the original AgeDB or DigiFace datasets as the true image set. Then, In all-shot evaluation, for AgeDB-IT2I-S and AgeDB-IT2I-M, we generate $100$ images per identity as the fake image set, and for AgeDB-IT2I-L, DigiFace-IT2I, VGGFace-IT2L and CIFAR-100-IT2I, we generate $20$ images per identity as the fake image set. In few-shot evaluation, we generate $500$ images per identity as the fake image set. For all generations, we employ the DDIM sampling technique [3] with $50$ steps. The prompt used during generation is “An image of {p}.” where “p” is the name of the identity (e.g., Albert Einstein).
@@
Appendix F Details for Implementation
@@
[…]
Appendix G Full Results
We report the performance of different methods in terms of FID score, human evaluation score, GPT-4o score, and DINO score in Table 7,  Table 8,  Table 9 and  Table 10, respectively. Across all tables, we observe that our PoGDiff consistently outperforms all baselines. More results and discussions are in  Appendix J.5. Notably, PoGDiff demonstrates significant improvements, especially in few-shot scenarios (i.e., for minority classes). It is also worth noting that CBDM [15] performs extremely poorly on AgeDB-IT2I-S and AgeDB-IT2I-M datasets. This is because their method samples text conditions from the entire space, which may work in one-hot class settings, but in our context (natural text conditions), this sampling technique misguides the model during training. In addition, Table 11 shows the gRecall for different methods on three datasets, AgeDB-IT2I-Small, AgeDB-IT2I-Medium, and AgeDB-IT2I-Large. These results show that our PoGDiff achieves much higher gRecall compared to all baselines, demonstrating its impressive diversity and coverage of different attributes of the same individual in the training set (see Appendix I for more discussion and examples on gRecall).
Table 7: Performance based on FID score.
Datasets (-IT2I) AgeDB DigiFace VGGFace CIFAR-100
Size Small Medium Large Large Large Large
Metric FID $\downarrow$
Shot All Few All Few All Few All Few All Few All Few
Vanilla 14.88 13.72 12.87 12.56 7.67 11.67 7.18 12.23 7.59 12.08 7.19 11.46
CBDM 14.72 14.13 11.63 11.59 7.18 11.12 6.96 12.72 7.23 11.91 7.26 11.94
T2H 14.85 13.66 12.79 12.52 7.61 11.64 7.14 12.22 7.34 12.02 7.10 11.39
PoGDiff (Ours) 14.15 12.88 10.89 10.64 6.03 10.16 6.84 11.21 6.29 10.97 6.24 9.41
Table 8: Performance based on DINO score.
Datasets (-IT2I) AgeDB DigiFace VGGFace CIFAR-100
Size Small Medium Large Large Large Large
Metric DINO (cosine similarity) scores $\uparrow$
Shot All Few All Few All Few All Few All Few All Few
Vanilla 0.42 0.37 0.39 0.28 0.34 0.25 0.42 0.36 0.41 0.29 0.48 0.32
CBDM 0.54 0.09 0.38 0.11 0.41 0.26 0.34 0.16 0.46 0.22 0.52 0.28
T2H 0.43 0.39 0.42 0.29 0.37 0.26 0.44 0.36 0.42 0.28 0345 0.30
PoGDiff (Ours) 0.77 0.73 0.69 0.56 0.66 0.52 0.64 0.49 0.69 0.55 0.73 0.61
Table 9: Performance based on human evaluation. The evaluation is a binary decision: Image is either judged as representing the same individual (score 1.0) or not (score 0.0).
Datasets (-IT2I) AgeDB VGGFace CIFAR-100
Size Small Medium Large Large Large
Metric Human Score $\uparrow$
Shot All Few All Few All Few All Few All Few
Vanilla 0.50 0.00 0.66 0.32 0.60 0.20 0.62 0.16 0.72 0.30
CBDM 0.50 0.00 0.44 0.08 0.56 0.12 0.54 0.10 0.63 0.24
T2H 0.50 0.00 0.66 0.32 0.60 0.20 0.62 0.16 0.72 0.30
PoGDiff (Ours) 1.00 1.00 0.96 0.92 0.84 0.68 0.78 0.64 0.84 0.72
Table 10: Performance based on GPT-4o evaluation. The scores are from 0 to 10, with higher scores indicating the individual resembles the well-known person.
Datasets (-IT2I) AgeDB VGGFace CIFAR-100
Size Small Medium Large Large Large
Metric GPT-4o Evaluation $\uparrow$
Shot All Few All Few All Few All Few All Few
Vanilla 5.20 3.20 4.30 2.90 4.90 3.60 4.50 2.90 6.00 3.20
CBDM 4.50 1.10 1.30 1.00 3.10 1.70 2.80 1.30 3.40 2.00
T2H 5.50 3.10 4.60 3.00 4.70 3.90 4.60 3.10 6.20 3.60
PoGDiff (Ours) 9.10 8.40 8.80 8.20 8.50 8.00 8.20 7.60 8.40 8.00
Datasets (-IT2I) AgeDB VGGFace CIFAR-100
Size Small Medium Large Large Large
Metric gRecall Score $\uparrow$
Shot All Few All Few All Few All Few All Few
Vanilla 0.017 0.000 0.104 0.167 0.196 0.200 0.133 0.167 0.200 0.160
CBDM 0.267 0.000 0.159 0.083 0.138 0.100 0.120 0.100 0.086 0.067
T2H 0.017 0.000 0.104 0.167 0.196 0.200 0.133 0.167 0.200 0.160
PoGDiff (Ours) 0.800 1.000 0.517 0.642 0.435 0.540 0.400 0.533 0.433 0.567
Appendix H Limitations
@@
Appendix I Additional Details for AgeDB-IT2I-small in Table 5
@@
Appendix J Discussion
J.1 Problem Settings
@@
Figure 6: Overview of label distributions for six IT2I datasets in bar plots. The x-axis corresponds to the identities (i.e., people or class).
Figure 7: TSNE visualization for all the methods for an example individual in the AgeDB-IT2M-large dataset.
J.2 More Detailed Analysis: Understanding Fig. 5
@@
J.3 Why Not Directly Smooth Text Embedding?
@@
J.4 Our Dataset Covers Different Levels of Sparsity